A table top is in the shape of a square. If the numerical values of twice the area of the square and thrice its perimeter added, the sum is 32. What is the measure of the side of the square table?
Answer by ikleyn(52778)

.

    2x^2 + 3*(4x) = 32


    2x^2 + 12x - 32 = 0


    x^2  +  6x  - 16 = 0


    (x+8)*(x-2) = 0.


The only positive root x= 2 is the solution.    ANSWER


CHECK.   2*2^2 + 3*(4*2) = 8 + 24 = 32.   ! Correct !

Solved.